# Weaver: Interweaving SQL and LLM for Table Reasoning

### v1

@@ Abstract
Querying tables with unstructured data is challenging due to the presence of text (or image), either embedded in the table or in external paragraphs, which traditional SQL struggles to process, especially for tasks requiring semantic reasoning. While Large Language Models (LLMs) excel at understanding context, they face limitations with long input sequences. Existing approaches that combine SQL and LLMs typically rely on rigid, predefined workflows, limiting their adaptability to complex queries. To address these issues, we introduce Weaver , a modular pipeline that dynamically integrates SQL and LLMs for table-based question answering (TableQA). Weaver generates a flexible, step-by-step plan that combines SQL for structured data retrieval with LLMs for semantic processing. By decomposing complex queries into manageable subtasks, Weaver improves accuracy and generalization. Our experiments show that Weaver consistently outperforms state-of-the-art methods across four TableQA datasets, reducing both API calls and error rates.
footnotetext: *These authors contributed equally to this work.   †Primary superviser of this work.
@@ 1 Introduction
Tables play a critical role across various domains such as finance (e.g., transaction records), healthcare (e.g., medical reports), and scientific research. However, many real-world tables are semi-structured Gupta et al. (2020), containing a combination of structured fields and unstructured content (such as free-form text or images), which makes reasoning and information retrieval challenging. Extracting insights from such data demands both logical and semantic reasoning. While SQL and Python-based methods excel in handling structured data, they fall short in dealing with unstructured text, missing entries, or implicit inter-column relationships.
Figure 1: Table and Question Answer example from WikiTQ Dataset
@@
Recent advances in Large Language Models (LLMs) have demonstrated strong capabilities in natural language understanding and contextual reasoning, opening new avenues for complex tasks. However, LLMs still face key limitations, particularly with long contexts and numerical or temporal reasoning. For instance, in the WikiTableQuestions Pasupat and Liang (2015) dataset, the query "Which country had the most competitors?" Figure 1 requires inferring the competitors’ countries from a "driver" column—information not explicitly present. While traditional tools like SQL or Python cannot resolve such gaps, LLMs can leverage their pre-trained knowledge to do so. Yet, the subsequent grouping and counting by country is something LLMs struggle with but SQL handles well. Solving such queries effectively requires a hybrid approach that combines the strengths of LLMs and programmatic methods. This raises a key question: Can SQL and LLMs be seamlessly interwoven?
@@
Some methods, such as Binder Cheng et al. (2022) and BlendSQL Glenn et al. (2024), integrate LLMs into SQL workflows by treating them as function calls. For example, Binder combines LLM reasoning with SQLite to support hybrid queries. While effective for simpler tasks, these approaches struggle with complex queries, as LLMs often fail to generate accurate multi-step logic. This highlights the need to decompose complex queries into smaller, manageable steps. Other approaches, like H-Star Abhyankar et al. (2024) and ReAcTable Zhang et al. (2024), use programmatic techniques to prune tables but rely heavily on costly API calls. Meanwhile, methods like ProTrix Wu and Feng (2024) limit reasoning to just two steps, making them insufficient for multi-hop queries. These rigid pipelines often constrain LLMs to final answer extraction and cannot handle questions requiring external or implicit knowledge.
@@
To address these challenges, we introduce a modular, planning-based framework that dynamically alternates between SQL for logical operations and LLMs for semantic reasoning. By decoupling these components, our approach overcomes the limitations of monolithic systems and significantly improves performance on complex Table QA tasks. The process begins with selecting relevant columns and generating natural-language descriptions to resolve schema inconsistencies. An LLM then generates a step-by-step reasoning plan, combining SQL queries for structured operations with LLM prompts for semantic inference or column augmentation. Each step produces an intermediate table, enabling transparent reasoning and easy backtracking. A final answer extraction step retrieves the result from the processed table. This flexible design integrates with standard database engines (e.g., MySQL, SQLite) and supports various deployment settings. Our approach offers the following key contributions:
@@
  • We propose Weaver , a modular and interpretable framework for hybrid query execution that dynamically decomposes complex queries into modality-specific steps (e.g., SQL, LLM, VLM) without manual effort.
@@
Figure 2: Step-by-step TableQA execution using Weaver approach.
@@ 2 Our Approach
This study addresses question answering tasks over tables containing both structured and unstructured data. Each instance consists of a table ($T$), a user query ($Q$), an optional paragraph ($P$), and a predicted answer ($A$). Queries span multiple categories: short-form queries (WikiTQ) involving direct lookups or aggregations; fact-checking queries (TabFact Chen et al. (2019)) that require claim verification; numerical reasoning queries (FinQA Chen et al. (2021b)) necessitating multi-step calculations; and multi-modal queries (FinQA and OTT-QA Chen et al. (2021a)) that demand reasoning over extensive textual contexts inside and outside tables. Complex queries, such as "Which country had the most competitors?", frequently require semantic inference when explicit data is unavailable. Previous approaches typically rely on single-step executions, limiting flexibility and interpretability. In contrast, our method dynamically integrates SQL for structured data operations and LLMs for semantic inference, providing adaptable and accurate query resolution.
@@ 2. Planning:
(a.) SQL Step: SQL steps manage structured data tasks, including filtering rows, formatting column data types, mathematical operations and data aggregations. For example, Figure 2 demonstrates an SQL step that generates an intermediate table, "unique_drivers."
(b.) LLM Step: LLM steps handle tasks beyond SQL’s capabilities, such as deriving new columns through semantic inference, sentiment analysis, or interpreting complex textual data. LLMs leverage either contextual paragraphs or their pretrained knowledge to perform these tasks. Each LLM step carefully integrates outputs back into the structured data tables to ensure coherence and consistency for subsequent SQL steps. Figure 2 illustrates how an LLM infers a "country" column from the "driver" column in the intermediate table "unique_drivers". The LLM is guided through structured prompts that leverage its pretrained knowledge and reasoning capabilities. Relevant information (extracted from external unstructured text in pre-processing step) is also passed to LLM if present.
Plan Verification.
Prompting techniques like self-refinement Madaan et al. (2023) and verification Weng et al. (2023) are known to enhance LLM reasoning by reducing errors and improving consistency. To leverage this, we use a secondary LLM to verify the initial plan, ensuring its logical consistency, robustness, and completeness. Gaps such as insufficient reasoning or formatting issues are addressed by refining the plan, as shown in step (D) in Figure 2. This verification improves the pipeline’s reliability and mitigates cascading errors.
@@ 3. Code Execution:
(a.) SQL Step - Query Generation: SQL operations involve formatting, filtering, joining, aggregating, and grouping data, with intermediate tables stored at each stage. SQL efficiently handles structured data operations, reducing reliance on LLM steps.
(b.) LLM Step - Prompt Generation: LLMs dynamically create prompts for operations exceeding SQL capabilities, such as generating new columns or interpreting textual insights. These LLM-generated prompts interact with structured intermediate tables from SQL steps, ensuring coherent integration.
@@ 4. Answer Extraction:
In the final pipeline stage, the intermediate table and user query are inputted to an LLM, which generates a natural language answer. Leveraging few-shot learning ensures output consistency and contextual accuracy, effectively resolving complex queries. Figure 2 illustrates this process with a sample TableQA example.
@@ 2.2 Optimization
We experimented to further enhance the efficiency of our pipeline, optimizing the planning strategy by minimizing unnecessary LLM calls and prioritizing SQL-based operations.
One key optimization involves pushing SQL operations such as filtering, aggregation, and formatting early in the pipeline as shown in Figure 3 in Appendix A.5. This reduces the volume of data that needs to be processed by the LLM, significantly reducing latency and computational overhead. Furthermore, parallelizing LLM inference across these optimized chunks further enhances efficiency, allowing the system to handle large-scale tabular reasoning tasks effectively, as shown Figure 3. To further streamline execution, sequential SQL steps are merged, reducing SQL calls and improving performance, see Figure 4 in Appendix A.5.
@@
Given the computational demands of LLMs for large tables we split data into smaller context-aware chunks before sending them to the LLM for inference. This prevents input truncation, maintains logical coherence between batches, and ensures optimal utilization of the context window of the model, as demonstrated in Figure 5, in the Appendix A.5. This optimization strategy further enhances Weaver planning and execution efficiency.
@@ 3 Experiments
Filtering Methodology. We define "hybrid" queries as those that require both SQL operations and LLM-based reasoning. These queries are more complex, as they necessitate not only structured data retrieval but also advanced reasoning capabilities, such as entity inference or free-text interpretation, which SQL alone cannot provide. To identify such queries, we use Binder-generated queries that incorporate user-defined LLM functions (UDFs). Queries involving UDFs indicate the need for semantic reasoning beyond SQL’s capabilities. It is important to note that we did not validate the correctness of Binder’s outputs; its role was purely to flag queries with hybrid characteristics. These candidate queries were then manually validated to confirm they genuinely required both SQL execution and LLM reasoning steps to arrive at the correct answer. For FinQA, we utilized the "qa" object to identify queries requiring multiple reasoning steps, excluding simple table lookups to ensure the selected queries involved more than just direct data retrieval. These flagged queries were then manually reviewed to confirm their need for both SQL execution and LLM reasoning.
@@
Dataset Statistics. After filtering, the hybrid versions of the datasets consist of: (a) WikiTQ: 510 examples (original: 4,344), (b) TabFact: 303 examples (original: 2,000), and (c) FinQA: 1,006 examples (original: 8,281). These represent the final queries filtered to create the hybrid versions: $\text{WikiTQ}_{\text{hybrid}}$, $\text{TabFact}_{\text{hybrid}}$, and $\text{FinQA}_{\text{hybrid}}$.
Evaluation Metrics
Traditional Exact Match requires an exact match between the model’s generated answer and the gold answer, which can unfairly penalize correct responses that differ only in format, for example, 2nd April 2024 and 04/02/2024 are semantically the same. To address this, we introduce Relaxed Exact Match (REM) metric which implements a three-step evaluation framework. First, we standardize the model’s output to align with the gold answer, handling variations such as unit representation and common abbreviations. For example, if the model returns the final table with column- "Year" with value- "17", and the expected answer is "17 years", we transform the output to match the reference format. Once the answers are format-aligned, we apply the standard EM metric to determine whether the processed output matches the gold answer. However, automated transformations can sometimes introduce unintended errors, such as incorrect unit conversions or context misalignment. To prevent such issues, we also perform human evaluation to ensure accurate answer matching.
LLM Models.
@@
In our research, we use state-of-the-art large language models (LLM) such as Gemini-2.0-Flash DeepMind (2024), GPT-4o-mini-2024-07-18, GPT-4o-2024-08-06 OpenAI et al. (2024) and the open-source DeepSeek R1-distill Llama-70B  [^1] https://github.com/meta-llama/llama- models/blob/main/models/llama3_3/LICENSE DeepSeek-AI et al. (2025), Shi et al. (2024) for table reasoning tasks. Our model inputs include in-context examples, the table, and the question for each step of the pipeline.
@@ 3.1 Baseline Methods
We evaluated our approach against several baselines that are broadly categorized into 4 categories: 1. Query Engines (Binder, BlendSQL): These methods generate hybrid SQL queries with LLMs as User Defined Functions. They leverage SQL to interpret tabular data and execute queries to retrieve relevant information., 2. End-to-End LLM QA: This approach leverages LLMs for question answering without intermediate query structuring. The model receives a query and table as input, generating answers based on learned patterns and reasoning. We employ GPT-4o, GPT-4o-mini, Gemini-2.0-Flash, and DeepSeek R1-distill LlaMA 70B for all tasks, 3. Pruning-Based Methods (ReacTable, H-Star): These methods first apply SQL or Python-based pruning techniques, such as filtering columns or rows, before passing the refined table to an LLM for final answer extraction, and 4. Planning Based Approach (ProTrix): ProTrix employs a two-step "Plan-then-Reason" framework. It first plans the reasoning, and assigns SQL to filter the table. Finally, it uses LLM to extract the final answer. By comparing our approach with these methods, we highlight the unique strengths of SQL-LLM Weaver , which combines SQL-based filtering with LLM-driven reasoning for more effective query resolution.
@@ 3.2 Results and Analysis
Our Weaver performs well on three challenging benchmarks; we present key findings next.
@@
The results in Table 1 compare GPT-3.5-turbo on the original dataset with GPT-4o-mini on the hybrid set. Despite leveraging a more capable model (GPT-4o-mini) for the hybrid queries, we observe substantial performance drops—H-Star and Binder see accuracy declines of 9.5% and 32.7%, respectively, on $\text{WikiTQ}_{\text{hybrid}}$.
@@
Notably, GPT-4o-mini outperforms GPT-3.5-turbo on benchmarks like MMLU and MATH (Source: OpenAI), yet still struggles on hybrid queries. This underscores their inherent difficulty and highlights the limitations of current methods in handling multi-step, semantically complex reasoning in Table QA.
@@
Table 2 demonstrates that Weaver consistently outperforms state-of-the-art baselines across all datasets. On WikiTQ, Weaver surpasses the best-performing baseline ProTrix by 5.5% across all four models. On TabFact, it achieves a breakthrough 91.2% using DeepSeek model, surpassing the 90% benchmark. On FinQA, it achieves accuracy of 65%, outperforming baselines by 4.6% on DeepSeek R1-distill Llama 70B.
WikiTQ TabFact FinQA
@@
DeepSeek R1-distill Llama 70B
@@
Weaver vs Query Engines. Binder and BlendSQL struggle with hybrid queries due to their rigid single-step execution framework. We observe that only 61% and 66% of the hybrid queries execute successfully in Binder and BlendSQL on WikiTQ. The reported accuracies for these methods are calculated based on successfully executed queries. Although BlendSQL slightly outperforms Binder with a modest 2% it frequently encounters type errors when integrating LLM-generated outputs into SQL operations. Weaver outperforms BlendSQL by 39%, 20.9% and 12.3% accuracy on WikiTQ, TabFact and FinQA using GPT-4o-mini.
Weaver vs Pruning Methods. H-Star and ReAcTable while effective for structured queries, perform poorly on semantic tasks. H-Star attains 59.0% and 63.5% accuracy on $\text{WikiTQ}_{\text{hybrid}}$ using GPT-4o-mini and Gemini-2.0-Flash, respectively, but struggles with row extraction. In some cases, its row-filtering heuristics discard critical contextual data essential for reasoning. H-Star’s higher accuracy on TabFact with GPT-4o stems from the dataset’s suitability for pruning techniques.  [^2] We didn’t test ReAcTable on FinQA since its prompts are tailored to other sets; modifying them changes baseline. Furthermore, Weaver  with GPT-4o-mini and DeepseekR1-Distill-LLAMA, despite their smaller size, performs competitively highlighting its effectiveness in resource-constrained environments where lightweight models are preferred.
Weaver vs Planning Method. ProTrix follows a two-step pipeline (planning and execution), achieves 61.4% and 62.2% on $\text{WikiTQ}_{\text{hybrid}}$ with GPT-4o-mini and Gemini-2.0-Flash. However, it fails in scenarios requiring intermediate semantic processing. For example, queries demanding dynamic column generation (e.g., inferring Country from Constructor) reveal its inability to seamlessly integrate SQL and LLM reasoning, leading to a 9% accuracy gap (GPT-4o) compared to Weaver .
@@
Gemini-2.0-Flash DeepSeek R1
@@
Table 3: Experimental results on short-form question answering on dataset with table and paragraphs.
@@
On ${}_{\text{MM}}$ dataset, Weaver achieved an accuracy of 53.02% using gpt-4o-mini model, significantly outperforming the end-to-end QA baseline, which scored 46.33%. These results highlight the strength of our modular, reasoning-driven approach, combining structured data (tables), unstructured data (text), and visual inputs (images) for superior performance. This underscores Weaver ’s versatility and scalability in addressing complex multimodal (table, text, images) QA tasks.
Efficacy Analysis.
We conducted an analysis to assess the effectiveness and scalability of Weaver in $\text{WikiTQ}_{\text{hybrid}}$, focusing on 98 large tables with over 30 rows and average token length of 17,731. Our results show that Weaver achieved an accuracy of 65.6%, outperforming ProTrix (37.5%) and H-Star (35.9%) by 28.1% and 29.7%, respectively, on these large tables Weaver maintained the same accuracy on the entire original dataset, demonstrating its ability to handle complex queries while remaining both scalable and robust. These results highlight Weaver ’s ability to tackle a wide range of table-based tasks with consistent performance.
In addition to delivering reliable performance, Weaver  offers the critical advantage of transparent, interpretable reasoning. By following a structured execution plan, it ensures that final answers are tightly aligned with preceding reasoning steps, enhancing traceability and reducing spurious outputs. This directly addresses a core limitation of large language models—hallucination and memorization. Unlike end-to-end LLMs, which may produce correct answers without valid reasoning, Weaver  only yields correct outputs when the underlying plan is sound, ensuring both reliability and interpretability.
Efficiency Analysis.
Table 4 demonstrate the efficiency of our proposed Weaver framework using number of API calls. We make six API calls which are much lower compared to approaches that use self-consistency (Binder) with 50 calls and H-Star which uses $\sim$ 8 calls to reach the answer.

### v2

@@ Abstract
Querying tables with unstructured data is challenging due to the presence of text (or image), either embedded in the table or in external paragraphs, which traditional SQL struggles to process, especially for tasks requiring semantic reasoning. While Large Language Models (LLMs) excel at understanding context, they face limitations with long input sequences. Existing approaches that combine SQL and LLM typically rely on rigid, predefined workflows, limiting their adaptability to complex queries. To address these issues, we introduce Weaver , a modular pipeline that dynamically integrates SQL and LLM for table-based question answering (Table QA). Weaver generates a flexible, step-by-step plan that combines SQL for structured data retrieval with LLMs for semantic processing. By decomposing complex queries into manageable subtasks, Weaver improves accuracy and generalization. Our experiments show that Weaver consistently outperforms state-of-the-art methods across four Table QA datasets, reducing both API calls and error rates.
footnotetext: *These authors contributed equally to this work.   †primary supervisor of this work.
@@ 1 Introduction
Tables play a critical role across various domains such as finance (e.g., transaction records), healthcare (e.g., medical reports), and scientific research. However, many real-world tables often contain a mix of structured fields alongside columns with embedded unstructured text (such as free-form text or images), which makes reasoning and information retrieval challenging. Extracting insights from such data demands both logical and semantic reasoning. While SQL and Python-based methods excel in handling structured data, they fall short in dealing with unstructured text, missing entries, or implicit inter-column relationships.
Figure 1: An exemplar hybrid query from the WikiTQ Dataset demonstrating the need for interwoven reasoning. The question “Which country had the most competitors?” requires the semantic inference of driver nationalities and the logical grouping and counting of competitors to produce the correct answer.
@@
Recent advances in Large Language Models (LLMs) have demonstrated strong capabilities in natural language understanding and contextual reasoning, opening new avenues for complex tasks. However, LLM still face key limitations, particularly with long contexts and numerical or temporal reasoning. For instance, in the WikiTableQuestions Pasupat and Liang (2015) dataset, the query “Which country had the most competitors?” Figure 1 requires inferring the competitors’ countries from a driver column information not explicitly present. While traditional tools such as SQL or Python cannot resolve such gaps, LLM can leverage their pre-trained knowledge to do so. Yet, the subsequent grouping and counting by country is something LLM struggle with but SQL handles well. Solving such queries effectively requires a hybrid approach that combines the strengths of LLM and programmatic methods. This raises a key question: Can SQL and LLM be seamlessly interwoven?
@@
Some methods, such as Binder Cheng et al. (2023) and BlendSQL Glenn et al. (2024), integrate LLM into SQL workflows by treating them as function calls. For example, Binder combines LLM reasoning with SQLite to support hybrid queries. While effective for simpler tasks, these approaches struggle with complex queries, as LLM often fail to generate accurate multi-step logic. This highlights the need to decompose complex queries into smaller, manageable steps. Other approaches, such as H-STAR Abhyankar et al. (2025) and ReAcTable Zhang et al. (2024), use programmatic techniques to prune tables but rely heavily on costly API calls. Meanwhile, methods like ProTrix Wu and Feng (2024) limit reasoning to just two steps, making them insufficient for multi-hop queries. These rigid pipelines often constrain LLM to answer extraction and cannot handle questions requiring external or implicit knowledge.
@@
To address these challenges, we introduce a modular, planning-based framework that dynamically alternates between SQL for logical operations and LLM for semantic reasoning. By decoupling these components, our approach overcomes the limitations of monolithic systems and significantly improves performance on complex Table QA tasks. The process begins with extracting relevant columns and generating column descriptions based on the given table and query, resolving formatting inconsistencies and ambiguities in table or column names. An LLM then generates a step-by-step reasoning plan, combining SQL queries for structured operations with LLM prompts for semantic inference or column augmentation. Each step produces an intermediate table, enabling transparent reasoning and easy backtracking. A final answer extraction step retrieves the result from the processed table. This adaptable design enables seamless integration with various database engines (e.g., MySQL, SQLite). Our approach offers the following key contributions:
@@
  • We propose Weaver , a modular and interpretable framework for hybrid query execution that dynamically decomposes complex queries into modality-specific steps (e.g., SQL, LLM, vision-language models (VLMs)) without manual effort.
@@
The code, along with other associated scripts, are available at https://coral-lab-asu.github.io/weaver.
Figure 2: The transparent and interpretable Weaver pipeline for Table QA. The framework dynamically alternates between SQL-based structured data operations (e.g., creating the unique_drivers table) and LLM-based semantic reasoning (e.g., inferring ‘country’ from driver names), with each step producing a traceable intermediate table to mitigate errors and enhance debugging.
@@ 2 Our Approach
This study addresses question answering tasks over tables containing both structured and unstructured data. Each instance consists of a table ($T$), a user query ($Q$), an optional paragraph ($P$), and a predicted answer ($A$). Queries span multiple categories: short-form queries (WikiTQ) involving direct lookups or aggregations; fact-checking queries (TabFact Chen et al. (2020)) that require claim verification; numerical reasoning queries (FinQA Chen et al. (2021b)) necessitating multi-step calculations; and multimodal queries (FinQA and OTT-QA Chen et al. (2021a)) that demand reasoning over extensive textual contexts inside and outside tables. Complex queries, such as “Which country had the most competitors?”, frequently require semantic inference when explicit data is unavailable. Previous approaches typically rely on single-step executions, limiting flexibility and interpretability. In contrast, our method dynamically integrates SQL for structured data operations and LLM for semantic inference, providing adaptable and accurate query resolution.
@@ 2. Planning:
(a.) SQL Step:
SQL steps manage structured data tasks, including filtering rows, formatting column data types, mathematical operations and data aggregations. For example, Figure 2 demonstrates an SQL step that generates an intermediate table, unique_drivers.
(b.) LLM Step:
LLM steps handle tasks beyond SQL’s capabilities, such as deriving new columns through semantic inference, sentiment analysis, or interpreting complex textual data. LLM leverage either contextual paragraphs or their pretrained knowledge to perform these tasks. Each LLM step carefully integrates outputs back into the structured data tables to ensure coherence and consistency for subsequent SQL steps. Figure 2 illustrates how an LLM infers a country column from the driver column in the intermediate table unique_drivers. The LLM is guided through structured prompts that leverage its pretrained knowledge and reasoning capabilities. Relevant information (extracted from external unstructured text in pre-processing step) is also passed to LLM if present.
Plan Verification:
Prompting techniques like self-refinement Madaan et al. (2023) and verification Weng et al. (2023) are known to enhance LLM reasoning by reducing errors and improving consistency. To leverage this, we use a secondary LLM to verify the initial plan, ensuring its logical consistency, robustness, and completeness. Gaps such as insufficient reasoning or formatting issues are addressed by refining the plan, as shown in planning step (D) (New Plan) in Figure 2. This verification improves the pipeline’s reliability and mitigates cascading errors.
@@ 3. Code Execution:
(a.) SQL Step - Query Generation:
SQL operations involve formatting, filtering, joining, aggregating, and grouping data, with intermediate tables stored at each stage. SQL efficiently handles structured data operations, reducing reliance on LLM steps.
(b.) LLM Step - Prompt Generation:
At each LLM step execution, we retrieve the inputs specified in the plan, namely the relevant column subset, source table, LLM prompt, and target column name. Using the prompt and selected columns, the LLM generates the target column values, which are then appended to the intermediate table.
@@ 4. Answer Extraction:
In the final pipeline stage, the intermediate table and user query are inputted to an LLM, which generates a natural language answer. Leveraging few-shot learning ensures output consistency and contextual accuracy, effectively resolving complex queries.
Figure 2 illustrates this process with a sample Table QA example.
@@ 2.2 Optimization
We experimented to enhance the efficiency of our pipeline, optimizing the planning strategy by minimizing unnecessary LLM calls and prioritizing SQL-based operations.
@@
One key optimization involves pushing SQL operations such as filtering, aggregation, and formatting early in the pipeline as shown in Figure 3. This reduces the volume of data that needs to be processed by the LLM, significantly reducing latency and computational overhead. Furthermore, parallelizing LLM inference across these optimized chunks further enhances efficiency, allowing the system to handle large-scale tabular reasoning tasks effectively, as shown in Figure 3. To further streamline execution, sequential SQL steps are merged, reducing SQL calls and improving performance, see Figure 4.
Given the computational demands of LLMs for large tables, we split data into smaller context-aware chunks before sending them to the LLM for inference. This prevents input truncation, maintains logical coherence between batches, and ensures optimal utilization of the context window of the model, as demonstrated in Figure 5, in the Appendix A.5. This optimization strategy further enhances Weaver  planning and execution efficiency.
@@ 3 Experiments
We define hybrid queries as those that require both SQL operations and LLM-based reasoning. These queries are more complex, as they necessitate not only structured data retrieval but also advanced reasoning capabilities, such as entity inference or free-text interpretation, which SQL alone cannot provide. To identify such queries, we use Binder-generated queries that incorporate user-defined LLM functions (UDFs). Queries involving UDFs indicate the need for semantic reasoning beyond SQL’s capabilities. We did not validate the correctness of Binder’s outputs; its role was purely to flag queries with hybrid characteristics. For FinQA, we utilized the ‘qa’ object to identify queries requiring multiple reasoning steps, excluding simple table lookups to ensure the selected queries involved more than just direct data retrieval. All candidate queries were then manually validated to confirm they require both SQL execution and LLM reasoning steps.
Dataset Statistics:
@@
After filtering, the hybrid versions of the datasets consist of: (a) WikiTQ: 510 examples (original: 4,344), (b) TabFact: 303 examples (original: 2,000), and (c) FinQA: 1,006 examples (original: 8,281). These represent the final queries filtered to create the hybrid versions: $\text{WikiTQ}_{\text{hybrid}}$, $\text{TabFact}_{\text{hybrid}}$, and $\text{FinQA}_{\text{hybrid}}$.
Evaluation Metrics:
Traditional Exact Match (EM) requires the model’s output to exactly match the gold answer, which can unfairly penalize semantically correct responses that differ only in formatting, such as 2nd April 2024 versus 04/02/2024. To address this limitation, we introduce the Relaxed Exact Match (REM) metric, which applies a three-step evaluation framework. First, we automatically standardize the model’s output format to align with the gold answer, handling differences such as units, date formats, casing, and common abbreviations. This normalization is performed using an LLM-based transformation prompt (see Appendix A), ensuring consistent resolution of trivial mismatches, for example, if the gold answer is 17 years and the model outputs 17, the unit is automatically appended to match the reference. Once both answers are format-aligned, we apply the standard EM metric to the normalized outputs, preserving the objectivity and reproducibility of exact string matching. Finally, in the rare cases where discrepancies remain, we incorporate a human safeguard to verify correctness and ensure that automatic normalization has not introduced errors or missed subtle equivalences. Importantly, this human evaluation is not a subjective judgment, but a strict check of semantic equivalence; for example, if the model outputs 5 feet 10 inches and the gold answer is 70 inches, the evaluator confirms whether the two represent the same quantity. By combining automatic normalization, objective exact matching, and a minimal safeguard, REM provides a fairer and more reliable evaluation compared to traditional EM.
LLM Models:
@@
In our research, we use state-of-the-art large language models (LLMs) such as Gemini-2.0-Flash DeepMind (2024), GPT-4o-mini-2024-07-18, GPT-4o-2024-08-06 OpenAI et al. (2024) and the open-source DeepSeek-R1-Distill-Llama-70B  [^1] https://github.com/meta-llama/llama-models/blob/main/models/llama3_3/LICENSE Guo et al. (2025), Shi et al. (2024) for table reasoning tasks. Our model inputs include in-context examples, the table, and the question for each step of the pipeline. We use the same LLM across all stages of the pipeline (planning, plan verification, code execution and final answer extraction), though the framework can be easily adapted to assign different LLMs to specific tasks. To ensure deterministic and stable output, we use a fixed temperature setting of 0.01.
@@ 3.1 Baseline Methods
We evaluated our approach against several baselines that are broadly categorized into 4 categories:
1. End-to-End LLM QA: This approach leverages LLM for question answering without intermediate query structuring. The model receives a query and table as input, generating answers based on learned patterns and reasoning. We employ GPT-4o, GPT-4o-mini, Gemini-2.0-Flash, and DeepSeek-R1-Distill-Llama-70B for all tasks.
2. Query Engines (Binder, BlendSQL): These methods generate hybrid SQL queries with LLM as User Defined Functions. They leverage SQL to interpret tabular data and execute queries to retrieve relevant information.
3. Pruning-Based Approach (ReacTable, H-STAR): These methods first apply SQL or Python-based pruning techniques, such as filtering columns or rows, before passing the refined table to an LLM for final answer extraction.
4. Planning-Based Approach (ProTrix): ProTrix employs a two-step “Plan-then-Reason” framework. It first plans the reasoning, and assigns SQL to filter the table. Finally, it uses LLM to extract the final answer. By comparing our approach with these methods, we highlight the unique strengths of Weaver , which combines SQL-based filtering with LLM-driven reasoning for more effective query resolution.
@@ 3.2 Results and Analysis
Weaver performs well on three challenging benchmarks, we present key findings next.
@@
The results in Table 1 compare GPT-3.5-turbo on the original dataset with GPT-4o-mini on the hybrid set. Despite leveraging a more capable model (GPT-4o-mini) for the hybrid queries, we observe substantial performance drops, H-STAR and Binder see accuracy declines of 9.5% and 32.7%, respectively, on $\text{WikiTQ}_{\text{hybrid}}$.
@@
Notably, GPT-4o-mini outperforms GPT-3.5-turbo on benchmarks like MMLU and MATH (Source: OpenAI  [^2] https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence), yet still struggles on hybrid queries. This underscores their inherent difficulty and highlights the limitations of current methods in handling multi-step, semantically complex reasoning in Table QA.
@@
Table 2 demonstrates that Weaver consistently outperforms state-of-the-art baselines across all datasets. On $\text{WikiTQ}_{\text{hybrid}}$, Weaver surpasses the best-performing baseline ProTrix by 5.5% across all four models. On $\text{TabFact}_{\text{hybrid}}$, it achieves a breakthrough 91.2% using DeepSeek model, surpassing the 90% benchmark. On $\text{FinQA}_{\text{hybrid}}$, it achieves accuracy of 65%, outperforming baselines by 4.6% on DeepSeek-R1-Distill-Llama-70B.
[…]
@@
DeepSeek-R1-Distill-Llama-70B
@@
Weaver vs Query Engines:
Binder and BlendSQL struggle with hybrid queries due to their rigid single-step execution framework. We observe that only 61% and 66% of the hybrid queries execute successfully in Binder and BlendSQL on $\text{WikiTQ}_{\text{hybrid}}$. The reported accuracies for these methods are calculated based on successfully executed queries. BlendSQL slightly outperforms Binder with a modest 2%. Weaver outperforms BlendSQL by 39%, 20.9% and 12.3% accuracy on $\text{WikiTQ}_{\text{hybrid}}$, $\text{TabFact}_{\text{hybrid}}$ and $\text{FinQA}_{\text{hybrid}}$ using GPT-4o-mini.
Weaver vs Pruning Methods:
H-STAR and ReAcTable while effective for structured queries, perform poorly on semantic tasks. H-STAR attains 59% and 63.5% accuracy on $\text{WikiTQ}_{\text{hybrid}}$ using GPT-4o-mini and Gemini-2.0-Flash, respectively, but struggles with row extraction. In some cases, its row-filtering heuristics discard critical contextual data essential for reasoning. H-STAR’s higher accuracy on $\text{TabFact}_{\text{hybrid}}$ with GPT-4o stems from the dataset’s suitability for pruning techniques.  [^3] We did not test ReAcTable on $\text{FinQA}_{\text{hybrid}}$ since its prompts are tailored to other sets; modifying them changes baseline. Furthermore, Weaver  with GPT-4o-mini and DeepSeek-R1-Distill-Llama-70B, despite their smaller size, performs competitively highlighting its effectiveness in resource-constrained environments where lightweight models are preferred.
Weaver vs Planning Method:
ProTrix follows a two-step pipeline (planning and execution), achieves 61.4% and 62.2% on $\text{WikiTQ}_{\text{hybrid}}$ with GPT-4o-mini and Gemini-2.0-Flash. However, it fails in scenarios requiring intermediate semantic processing. For example, queries demanding dynamic column generation (e.g., inferring Country from driver name) reveal its inability to seamlessly integrate SQL and LLM reasoning, leading to a 9% accuracy gap (GPT-4o) compared to Weaver .
Weaver on Multimodal Data:
@@
Gemini-2.0-Flash DeepSeek-R1
@@
Table 3: Experimental results on multimodal QA (tables + paragraphs). : DeepSeek-R1 refers to DeepSeek-R1-Distill-Llama-70B.
@@
[…]
Efficacy Analysis:
We conducted an analysis to assess the effectiveness and scalability of Weaver in $\text{WikiTQ}_{\text{hybrid}}$, focusing on 98 large tables with over 30 rows and average token length of 17,731 per table. Our results show that Weaver achieved an accuracy of 65.6%, outperforming ProTrix (37.5%) and H-STAR (35.9%) by 28.1% and 29.7%, respectively. On these large tables, Weaver achieved 65.6%, the same accuracy as on the entire dataset, demonstrating its ability to handle complex queries while remaining both scalable and robust. These results highlight Weaver ’s ability to tackle a wide range of table-based tasks with consistent performance.
In addition to delivering reliable performance, Weaver  offers the critical advantage of transparent, interpretable reasoning. By following a structured execution plan, it ensures that final answers are tightly aligned with preceding reasoning steps, enhancing traceability and reducing spurious outputs. This directly addresses a core limitation of large language models, hallucination and memorization. Unlike end-to-end LLM, which may produce correct answers without valid reasoning, Weaver  only yields correct outputs when the underlying plan is sound, ensuring both reliability and interpretability.
Efficiency Analysis:
Table 4 demonstrate the efficiency of our proposed Weaver framework using number of API calls. We make about six API calls which are much lower compared to approaches that use self-consistency (Binder) with 50 calls and H-STAR which uses $\sim$ 8 calls to reach the answer.